Two blocks on a frictionless table are held at rest with a compressed spring between them. When the blocks are released, the spring pushes them apart and they move in opposite directions. The spring has a spring constant of 215N / m and was compressed by 0.20 m. If one of the blocks moves at 1.50 m/s and the other moves away at 3.00 m/s , what is the mass of the heavier block in kg? (Hint: mechanical energy in this system is conserved)
Answer #1

Let m1 and m2 are masses of blocks . Let m1 be heavier than m2

By Conservation of energy , (1/2) k x2 = (1/2) m1 v12 + (1/2) m2 v22 ..........(1)

Where k = 215 N/m is spring constant , x = 0.2 m is compressed distance of spring , v1 is speed of block of mass m1 , v2 is speed of block of mass m2

By Conservation of momentum , m1 v1 - m2 v2 = 0

( LHS is total momentum immediately after blocks are released and RHS is initial momentum    which is zero ) .

hence we get , m1 = ( v2 / v1 ) m2 = ( 3 / 1.5 ) m2 = 2 m2 ..............(2)

Using eqn.(2) and by substituting values of k and x , we rewrite eqn.(1) as

8.6 = m2 { ( 2 \times 1.5 \times 1.5 ) + ( 3 \times 3 ) }   or    m2 = 0.637 kg

Heavier mass m1 = 2 m2 = 2 \times 0.637 = 1.274 kg